Circular No. 7139

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

MXB 1659–29 C. B. Markwardt, National Research Council and Goddard Space

Flight Center (GSFC); E. Smith and J. Swank, GSFC; and J. in ’t Zand, Space Research Organization of the Netherlands, Utrecht, write: “Observations of the transient source MXB 1659–29 (IAUC 7138) were performed by RXTE on Apr. 5.83–6.05 UT. The source was detected at a level of ∼ 38 mCrab (2–10 keV), but was punctuated by a series of x-ray dips that lasted 5–350 s. A single x-ray eclipse was observed between Apr. 5.97206 and 5.98257 (barycenter-corrected). Cominsky and Wood (1989, Ap.J. 337,

485) have published an orbital period of 7.1161138(17) hr, based on previous activity in 1976–1978, and we conclude that the observed eclipse is number 27708 according to their numbering convention. The observed center of the eclipse arrived 7.0 ± 2.6 min before the predicted time, based on a constant period, implying an orbital Ṗ = (−4.8 ± 1.8) × 10^{−11}. No xray bursts were seen, and no pulsations or QPOs were detected. Follow-up observations are being performed by RXTE.”

HERCULES X-1 A. M. Levine, Center for Space Research, Massachusetts Institute of

Technology (MIT); and R. Corbet, Goddard Space Flight Center (GSFC), NASA, and Universities Space Research Association, report on behalf of the RXTE/ASM team at MIT and NASA/GSFC: ”The 35-day longterm cycle has been obvious in the RXTE All-Sky Monitor 1.5–12-keV lightcurve for Hercules X-1, with the ‘main high state’ (see, e.g., Scott and Leahy 1999, Ap.J. 510, 974) being clearly detected every 35 days for > 3 yr. The first exception to this regular appearance of the main high states has recently occurred. No main high state appeared as expected around Mar. 23, with an upper limit of ∼ 15 mCrab. Earlier main high states typically reached peak intensities of 40–100 mCrab. This may represent an expansion of the accretion disk, such that the neutron star remains occulted through the entire 35-day cycle. Optical observations to determine whether the companion, HZ Her, is still being heated by x-rays are encouraged.”

SUPERNOVA 1999aa IN NGC 2595 CCD magnitude estimates, unfiltered unless otherwise noted: Mar.

13.666 UT, 16.5 (S. Yoshida and K. Kadota, Ageo City, Japan); 16.534, 16.2 (Yoshida); 18.826, V = 15.6 ± 0.2 (L. Kiss and B. Csák, Piszkéstető, Hungary); 20.92, 16.0 (S. Moretti and S. Tomaselli, Forlí, Italy).

1999 April 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green